# MANGO: Multimodal Attention-based Normalizing Flow Approach to Fusion Learning

### v1

@@ 1 Introduction
Human perceptions interpret the surrounding world in a multimodal way via multiple input channels, such as vision, text, or audio. The deep learning-based multimodal fusion methods have majorly improved the performance of various problems, e.g., classification [20, 38, 37, 58], action recognition [12, 54, 57], semantic segmentation [65, 56, 59, 61, 60], object detection [72]. The recent large multimodal models, e.g., ChatGPT [1], Gemini [52], Chaemelon [51], LLaMMA [53], etc, introduced for general-assistant purposes have also shown impressive performance on these applications.
@@ 2.2 Limitations of Normalizing Flows
$\scriptsize\begin{split}\mathbf{X}_{1},\mathbf{X}_{2}&=\operatorname{partition}(\mathbf{X})\\ \mathbf{Y}_{1}&=\mathbf{X}_{1},\quad\mathbf{Y}_{2}=\mathbf{X}_{2}\odot\exp\left(\mathcal{S}(\mathbf{X}_{1})\right)+\mathcal{T}(\mathbf{X}_{1})\\ \mathbf{Y}&=\operatorname{merge}([\mathbf{Y}_{1},\mathbf{Y}_{2}])\end{split}$ (1)
@@ 3.1 The Proposed Invertible Cross-Attention (ICA)
$\scriptsize\begin{split}\mathbf{X}_{1},\mathbf{X}_{2}&=\operatorname{partition}([\mathbf{x}_{1},...,\mathbf{x}_{N}])\\ \mathbf{Q}&=\operatorname{LN}(\operatorname{LP}(\mathbf{X}_{1})),\quad\mathbf{K}=\operatorname{LN}(\operatorname{LP}(\mathbf{X}_{1})),\quad\mathbf{V}=\mathbf{X}_{2}\\ \mathbf{Y}_{1}&=\mathbf{X}_{1},\quad\mathbf{Y}_{2}=\operatorname{softmax}\left(\frac{\mathbf{Q}\times\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V}\\ \mathbf{Y}&=\operatorname{merge}([\mathbf{Y}_{1},\mathbf{Y}_{2}])\end{split}$ (3)
@@
$\scriptsize\begin{split}\mathbf{Y}_{1},\mathbf{Y}_{2}&=\operatorname{partition}([\mathbf{y}_{1},...,\mathbf{y}_{N}])\\ \mathbf{Q}&=\operatorname{LN}(\operatorname{LP}(\mathbf{Y}_{1})),\quad\mathbf{K}=\operatorname{LN}(\operatorname{LP}(\mathbf{Y}_{1})),\quad\mathbf{V}=\mathbf{Y}_{2}\\ \mathbf{X}_{1}&=\mathbf{X}_{1},\quad\mathbf{X}_{2}=\left[\operatorname{softmax}\left(\frac{\mathbf{Q}\times\mathbf{K}^{T}}{\sqrt{d}}\mathbf{M}\right)\right]^{-1}\mathbf{V}\\ \mathbf{X}&=\operatorname{merge}([\mathbf{X}_{1},\mathbf{X}_{2}])\end{split}$ (5)

### v2

@@ 1 Introduction
Human perceptions interpret the surrounding world in a multimodal way via multiple input channels, such as vision, text, or audio. The deep learning-based multimodal fusion methods have majorly improved the performance of various problems, e.g., classification [20, 42, 37, 67, 62, 42], action recognition [12, 61, 66], semantic segmentation [74, 65, 68, 64, 70, 69, 70], object detection [81, 86, 41], scene graph generation [40, 39], human-brain understanding [43, 45, 44]. The recent large multimodal models [38, 27, 28], e.g., ChatGPT [1], Gemini [59], Chaemelon [58], LLaMA [60], etc, introduced for general-assistant purposes have also shown impressive performance on these applications.
@@ 2.2 Limitations of Normalizing Flows
$\scriptsize\begin{split}\mathbf{X}_{1},\mathbf{X}_{2}&=\operatorname{partition}(\mathbf{X})\\ \mathbf{Y}_{1}&=\mathbf{X}_{1},\qquad\mathbf{Y}_{2}=\mathbf{X}_{2}\odot\exp\left(\mathcal{S}(\mathbf{X}_{1})\right)+\mathcal{T}(\mathbf{X}_{1})\\ \mathbf{Y}&=\operatorname{merge}([\mathbf{Y}_{1},\mathbf{Y}_{2}])\end{split}$ (1)
@@ 3.1 The Proposed Invertible Cross-Attention (ICA)
$\scriptsize\begin{split}\mathbf{X}_{1},\mathbf{X}_{2}&=\operatorname{partition}([\mathbf{x}_{1},...,\mathbf{x}_{N}])\\ \mathbf{Q}&=\operatorname{LN}(\operatorname{LP}(\mathbf{X}_{1})),\qquad\mathbf{K}=\operatorname{LN}(\operatorname{LP}(\mathbf{X}_{1})),\qquad\mathbf{V}=\mathbf{X}_{2}\\ \mathbf{Y}_{1}&=\mathbf{X}_{1},\qquad\mathbf{Y}_{2}=\operatorname{softmax}\left(\frac{\mathbf{Q}\times\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V}\\ \mathbf{Y}&=\operatorname{merge}([\mathbf{Y}_{1},\mathbf{Y}_{2}])\end{split}$ (3)
@@
$\scriptsize\begin{split}\mathbf{Y}_{1},\mathbf{Y}_{2}&=\operatorname{partition}([\mathbf{y}_{1},...,\mathbf{y}_{N}])\\ \mathbf{Q}&=\operatorname{LN}(\operatorname{LP}(\mathbf{Y}_{1})),\qquad\mathbf{K}=\operatorname{LN}(\operatorname{LP}(\mathbf{Y}_{1})),\qquad\mathbf{V}=\mathbf{Y}_{2}\\ \mathbf{X}_{1}&=\mathbf{X}_{1},\qquad\mathbf{X}_{2}=\left[\operatorname{softmax}\left(\frac{\mathbf{Q}\times\mathbf{K}^{T}}{\sqrt{d}}\mathbf{M}\right)\right]^{-1}\mathbf{V}\\ \mathbf{X}&=\operatorname{merge}([\mathbf{X}_{1},\mathbf{X}_{2}])\end{split}$ (5)
@@ 5 Conclusions
Acknowledgment. This work is partly supported by NSF CAREER (No. 2442295), NSF SCH (No. 2501021), NSF E-RISE (No. 2445877), NSF SBIR Phase 2 (No. 2247237) and USDA/NIFA Award. We also acknowledge the Arkansas High-Performance Computing Center (HPC) for GPU servers. Nitin Agarwal’s participation was supported by U.S. NSF (OIA-1946391, OIA-1920920), AFOSR (FA9550-22-1-0332), ARO (W911NF-23-1-0011, W911NF-24-1-0078, W911NF-25-1-0147), ONR (N00014-21-1-2121, N00014-21-1-2765, N00014-22-1-2318), AFRL, DARPA, Australian DSTO Strategic Policy Grants Program, Arkansas Research Alliance, the Jerry L. Maulden/Entergy Endowment, and the Donaghey Foundation at the University of Arkansas at Little Rock.
Appendices